OpenStudy (anonymous):

simplify: f(2) f(-3) f(k) f(kpower2-1)=

OpenStudy (anonymous):

This isnt clear for me.

OpenStudy (anonymous):

problem: f(x)=2/x-3

OpenStudy (anonymous):

ok just put in the values into the equation. so for example f(2)= 2/2 -3= -2 (or is it 2/(2-3)?)

OpenStudy (anonymous):

for k u do the same, so f(k)=2/k-3 f(k^2-1)= 2/(k^2-1)-3